John is twice as old as Sarah. Three years from today, the sum of their ages will be 42. How old is John today?

Answer by robertb(5830)

j = 2s,
j+3 + s+3 = 42, or j + s = 36. Substituting the 1st equation into the last equation, we get 2s + s = 36, or 3s = 36, or s = 12, the age of Sarah now, and j = 2s = 24, the age of John.